OpenStudy (anonymous):

Factor x^3 + 2x^2 9x - 18, given that -2 is a zero.

OpenStudy (marissalovescats):

You could do synthetic division and then it will be facorable

OpenStudy (anonymous):

Option A: (x 2)(x2 9) = 0 Option B: (x + 2)(x 3)(x + 3) = 0 Option C: (x 2)(x 3)(x + 3) = 0 Option D: \[(x+2)(x-\sqrt{3})(x+\sqrt{3})=0\]

OpenStudy (marissalovescats):

x^2+2x^2+9x-19 right?

OpenStudy (anonymous):

yes

OpenStudy (marissalovescats):

That one can't be solved by synthetic You'll know that the factor will have x+2 in it because -2 is a solution

OpenStudy (anonymous):

So that eliminates A and C

OpenStudy (marissalovescats):

And we know the first one can't be it because of the x^3, you'll have to multiply is by x twice and get x^5 which doesn't fit in x^3 + 2x^2 9x - 18

OpenStudy (marissalovescats):

So that leaves D

OpenStudy (marissalovescats):

Choice B does say: (x+2)(x^3)(x+3) Right?

OpenStudy (anonymous):

no it says (x+2)(x-3)(x+3)-0 my minus signs stick and don't always appear

OpenStudy (marissalovescats):

Oh then yes B is the answer

OpenStudy (marissalovescats):